# The Design Space of Lockfiles Across Package Managers

### v2

@@ 4.3.4 Debugging benefit
Almost all developers have experienced breaking dependency updates and semantic versioning (semver) violations in their projects. Developers who commit lockfiles use them to help in debugging breaking changes. P2 recollect how they copied a previous version of a lockfile to resolve a breaking update, ”That was one time I literally manually copy-pasted. I went exactly to where that package was resolved before, and it was working, and I literally replaced it”.
Another advantage of using locked versions is that developers can explicitly control when to allow dependency updates, preventing unexpected breakages. P11 mention that they update the lockfile only when they are prepared to handle potential breaking changes. (“From time to time, I run poetry updates with the intention of getting the latest and being in a mindset where I now have the surplus time and energy to handle breaking changes or issues that could arise from using the latest versions.”) P7 mention how they have to enforce the lockfile to prevent a dependency from getting updated because it does not follow the semantic versioning guidelines. (”So basically, it was the same problem. Semantic versioning was not respected by them. So I updated the README, and saying that if they download from cargo, they should use cargo install –locked. Because in that way, I’m sure that they’re using the dependency in the lockfile.”).
In Cargo, there is the yanked release issue, when dependency is no longer available by version, but remain available by checksum. The Cargo dependency resolver ignores yanked versions unless they are specified in the lockfile, occasionally causing the build to fail. P6 discuss how they overcome the yanked version problem in Cargo through lockfiles, ”If you have the lockfile, you can always build it, even if it’s yanked. But if you don’t have a lockfile and it gets yanked, you cannot under any circumstances build it.”.
@@ 4.3.5 Security benefit
Another key point raised by the developers is that lockfiles are indirectly used by other security scanning tools. For instance, P1 mention using the Black Duck software component analysis tool, which relies on lockfiles (”Black Duck uses a list of vulnerabilities based on the package names, on their versions, and reports on a regular basis on our products to know if one of the dependencies of this product is having a known vulnerability. So that’s one thing which leverages the package-lock.json”). When a lockfile is present, dependency management bots such as Dependabot can also detect and flag vulnerabilities within indirect dependencies (”It’s definitely beneficial to have a lockfile for the purposes of Dependabot. Yeah. Dependabot is what I rely on heavily for making sure I’m dealing with supply chain security properly… So Dependabot will scan the lockfile in order to determine the whole dependency tree, and then it wll warn you for CVEs even within transitive dependencies”, P10).
@@ 4.4 Results for RQ4: What challenges do developers face when managing lockfiles?
As part of our interview protocol, we ask developers who do not commit lockfiles about their reasons, and those who do commit lockfiles about the challenges they face when using them. From their responses, we identify several common challenges, including transitive locking with libraries, managing dependency updates, limited readability, cache invalidation, and the learning curve associated with lockfile practices.
4.4.1 Indirect dependency locking challenge
The main reason mentioned by four out of six developers who do not commit lockfiles is that their project is a library rather than an application or executable. These projects are in Python, Javascript or Rust and are in various application domains. The challenge they face is that even if they lock the versions of dependencies in their project, they cannot publish a lockfile to the package registry, and downstream dependents could not enforce the lockfile for their library. As a result, they choose not to enforce a lockfile in their development environment, as doing so would give them a false sense of security. (”So basically it is better not to commit a lockfile because then you kind of run into these issues faster. And, uh, yeah, you don’t have this false sense of security where you think it’s all locked, but in reality, everybody who uses it doesn’t lock it.”, P4). Following the same reasoning, P8 explain that they prioritize the experience of the users of their library over their own developer experience, and as a result, they do not commit lockfiles.
Similarly, some developers view determinism as a concern of downstream users who depend on their library. (”You could make your end thing deterministic, but as a library author, for me, it’s, that’s not really my problem.” P12).
P9 do not currently commit lockfiles and they mention that they would reconsider their decision based on the updated Cargo guidelines introduced in 2023. They refer to the StackOverflow thread Should Cargo.lock be committed when the crate is both a Rust library and an executable. (”This was the previous recommendation on StackOverflow, but has since changed. I have not revisited this choice, but if I were to start a new project today, I would follow the latest recommendations and defaults”, P9.)
@@ 4.4.2 Slowing down dependency updates
P11 mention that, with Poetry, if a lockfile references a version of a dependency that has been yanked, it results in an error, unlike Cargo. However, if they do not have a lockfile, Poetry will resolve a newer version, without raising an error. They explain the issue with an example. ”Now, if you, at 9 a.m. in the morning, let’s say a library pushes a new version. At 9.30, you create a lockfile with that version locked. And at 10 a.m., they delete that version from PyPI. Well, then anybody who syncs using that lockfile hits a big issue. Because that version is no longer available”. At the same time, they acknowledge that this behavior can sometimes be helpful. ”it reports it as yanked, and therefore doesn’t install it. So you get an early error, which in some situations is desirable”.
In npm, the current state of project dependencies is referred to as the ideal state. For projects without a lockfile, npm CLI always resolves dependencies to this ideal state at the time of installation. P5 prefer resolving this ideal state rather than relying on a potentially outdated lockfile. ”Because you’re always resolving the ideal tree, like at a given moment, that tends to be a better resolution than whatever state your lockfile is accumulated over six months or whatever”. They also believe that continuously resolving the ideal tree helps fixing bugs and vulnerabilities more effectively. ”Oftentimes, either you’ll just miss the defect because it will be introduced for a short period of time and then got published with a fix. And so you just won’t even notice it just because you’re always resolving the ideal tree”.
@@ 4.4.3 Human readability challenge
All five npm/pnpm developers mention that they find the lockfiles too long, verbose and hard to read. For example, P1 state ”it’s quite wild what you can find inside the package lock. So it’s not really readable”, whereas P2 mention how this verbosity can complicate the code reviews and continuous integration, ”Git is kind of overloading and taking up a bunch of memory. And is even having trouble showing them all”. P5 also comment on the human readability of lockfiles, ”save us from the insanity that these machine generated lock files, which are just, you know, they’re human readable, but they’re not because they’re so big”. Poetry/Pipenv developers raise the same issue where P10 explain, ”it’s like a huge JSON file. So it’s kind of annoying to review the diff. And I don’t get a lot of value out of reviewing the diff, necessarily.”. P11 add that they also find it difficult to code review the lockfile, ”I will be much more likely to review any changes to a pyproject.toml or, you know, direct dependency file. So I’d say it’s to some degree a weakness of lockfiles that they are so big”.
@@
On the other hand, none of the Cargo or Go developers raise issues related to the readability. Go developer P15 specifically mention they find reviewing go.mod files very easy. ”But usually my Go mod files are very like, they do not go over one page of my screen”. P12 who develop in Python and do not use lockfiles also comment that they get a better developer experience with Go, ”It’s basically almost the same because go.mod is also kind of very minimalistic, So I guess you get all of the benefits of like, you know, pip lockfile or poetry lockfile or whatever with go without having to think about it.” This confirms the fact that, depending on the lockfile implementation details, such as the content included in lockfiles, developers either effectively use them for manual code reviews or tend to overlook them entirely.
@@
When asked about the information they consider during code reviews of lockfiles, the majority of developers who commit lockfiles mention that they focus on the name, URL, version of the package, and the number of indirect dependencies it includes. A consistent observation we make is that none of the developers pay attention to the checksum of the dependencies in the lockfile. (I’ve never actually checked checksums, P7). Following the same reasoning, Go developers recommend the Go’s approach of having checksums in a separate file. (”To be honest, I really don’t care where the checksum is stored. It makes it clean to have it separated because I only have to put it into the Go mod file and then the checksum will be generated automatically in the sum file. So that’s like having a look at your Go mod file is much cleaner., P15 ”).
@@ 4.4.4 Cache challenge
A common issue faced by all developers is cache invalidation. P8 remark that ”Cargo caching mechanisms are buggy,” whereas P1 mention that they have to recreate the lockfile to address cache invalidation problems, ”I would say most of the time what people do is they delete the node modules folder, they delete the package lock, and then they retry and install”. Following up on cache invalidation, P2 mention that they occasionally run into operating system specific issues which cannot be solved by simply deleting the cache (”Just working across different operating systems, um, you know, also on macOS. I’ve had issues where deleting, like doing that same process, deleting the node modules folder, created issues with npm and then nothing would install.”). Go developers also face cache related issues with Go proxy when packages get re-released under the same version (”Go proxy already had that basically broken release downloaded and stored in its proxy cache. So people were like opening issues that they had checksum issues”, P15).
@@ 4.4.5 Learning challenge
Another point raised by developers is the learning curve of lockfile related practices across different package managers. P14 compare Gradle and Go noting that they find it easier to learn and understand dependency resolution practices in Go, ”it’s very, very simple. It doesn’t have much of a learning curve, such as like Maven or, Gradle”. P12 also comment on this ”I think Go is just an easier, it’s just easier. It works better. And, and, you know, it’s a newer language and lessons learned”.
npm developers often mention that the logic behind dependency resolution and lockfile creation processes are unclear (”maybe an understanding of how it constructs the lockfile. Cause it’s still kind of a mystery to me”, P2). Many also comment on the unhelpfulness of error messages when debugging lockfile related issues (”The error message that you get is really cryptic.”, P1). They note that resolving these issues requires an in-depth understanding of the underlying processes, and the debugging itself is extremely time consuming (”Some of these issues I’ve had with it are just so hard to debug until I do some very thorough digging of what’s going on”, P2). P15 compare npm CLI with Go and explain that the minimalism makes the dependency resolution process in Go more transparent and easier to understand. ”You just say npm install and like a wall of text goes down and you have a gazillion packages installed and you don’t really know what happens and why those are installed. Compare that to Go, which is, in my opinion, very open and very clear on what dependencies you are using.
@@ 6.1 Dependency resolution and package managers
One of the key responsibilities of a package manager is resolving compatible versions for all the dependencies in a project. Zacchiroli et al.[19] review the usage of dependency resolving algorithms, including specialized constraint solving algorithms as well as ad-hoc dependency tree based algorithms. They conduct a census of dependency resolution mechanisms across 18 package managers, using version locking as one of the criteria for comparison. Pinckney et al.introduce a unified framework built on top of Max-SMT solvers to resolve dependencies more systematically, moving beyond ad-hoc algorithms [20].
@@
Liu et al.[10] analyze vulnerability propagation within dependency trees by applying npm-specific dependency resolution rules, rather than relying on simple reachability analysis for indirect dependencies. They recommend managing dependencies with lockfiles. Along similar lines, He et al.[27] study the effect of dependency pinning in npm projects. They find that local pinning leads to more security vulnerabilities due to bloated and outdated dependencies. They also suggest that the risk of malicious package updates can be reduced when a small set of core dependencies pin their versions and keep them updated regularly.
@@ 6.3 Adoption of dependency management tools within development teams
Previous works have also used combined qualitative and quantitative methods to learn about developers’ adoption of dependency management tools. Schorlemmer et al.[38] study the adoption of software signing in four package registries; Maven, PyPI, DockerHub and Huggingface. Kalu et al.[39] conduct a follow-up interview study to understand software signing in practice. They identify that strict signature rules increase the quantity of signatures, and that registry policies have an impact on the developer decision to adopt proper signing.
Saroar et al.[40] survey 90 developers to understand how they use, choose, and troubleshoot GitHub Actions. Holtgrave et al.[41] analyze different attack types that can be performed by impersonating GitHub user accounts. Bifolco et al.[42] assess the accuracy of the GitHub dependency graph in Java and Python projects. Bifolco et al.use lockfiles as one source of truth to assess the GitHub dependency graph.
Dependency management bots such as dependabot, renovate and Greenkeeper automate the process of updating dependency specification files and lockfiles when new dependency versions are available. Alfadel et al.[43] and Mohayeji et al.[44] evaluate how developers respond to security updates suggested by dependabot. Rombaut et al.[45] study whether Greenkeeper reduces developer effort or introduces unnecessary workload. They mention lockfiles as a way to overcome in-range breaking changes, and also briefly mention about the challenge library developers face for sharing lockfiles. Our qualitative analysis confirms that developers do use bots to maintain dependency specification files and lockfiles up to date.

### v3

@@ 4.3.4 Debugging benefit
Almost all developers have experienced breaking dependency updates and semantic versioning (semver) violations in their projects. Developers who commit lockfiles use them to help in debugging breaking changes. P2 recollect how they copied a previous version of a lockfile to resolve a breaking update, “That was one time I literally manually copy-pasted. I went exactly to where that package was resolved before, and it was working, and I literally replaced it”.
Another advantage of using locked versions is that developers can explicitly control when to allow dependency updates, preventing unexpected breakages. P11 mention that they update the lockfile only when they are prepared to handle potential breaking changes. (“From time to time, I run poetry updates with the intention of getting the latest and being in a mindset where I now have the surplus time and energy to handle breaking changes or issues that could arise from using the latest versions.”) P7 mention how they have to enforce the lockfile to prevent a dependency from getting updated because it does not follow the semantic versioning guidelines. (“So basically, it was the same problem. Semantic versioning was not respected by them. So I updated the README, and saying that if they download from cargo, they should use cargo install –locked. Because in that way, I’m sure that they’re using the dependency in the lockfile.”).
In Cargo, there is the yanked release issue, when dependency is no longer available by version, but remain available by checksum. The Cargo dependency resolver ignores yanked versions unless they are specified in the lockfile, occasionally causing the build to fail. P6 discuss how they overcome the yanked version problem in Cargo through lockfiles, “If you have the lockfile, you can always build it, even if it’s yanked. But if you don’t have a lockfile and it gets yanked, you cannot under any circumstances build it”.
@@ 4.3.5 Security benefit
Another key point raised by the developers is that lockfiles are indirectly used by other security scanning tools. For instance, P1 mention using the Black Duck software component analysis tool, which relies on lockfiles (“Black Duck uses a list of vulnerabilities based on the package names, on their versions, and reports on a regular basis on our products to know if one of the dependencies of this product is having a known vulnerability. So that’s one thing which leverages the package-lock.json”). When a lockfile is present, dependency management bots such as Dependabot can also detect and flag vulnerabilities within indirect dependencies (“It’s definitely beneficial to have a lockfile for the purposes of Dependabot. Yeah. Dependabot is what I rely on heavily for making sure I’m dealing with supply chain security properly… So Dependabot will scan the lockfile in order to determine the whole dependency tree, and then it wll warn you for CVEs even within transitive dependencies”, P10).
@@ 4.4 Results for RQ4: What challenges do developers face when managing lockfiles?
As part of our interview protocol, we ask developers who do not commit lockfiles about their reasons, and those who do commit lockfiles about the challenges they face when using them. From their responses, we identify several common challenges, including indirect locking with libraries, managing dependency updates, limited readability, cache invalidation, and the learning curve associated with lockfile practices.
4.4.1 The challenge of locking library dependencies
The main reason mentioned by four out of six developers who do not commit lockfiles is that their project is a library rather than an application or executable. These projects are in Python, Javascript or Rust and are in various application domains. The challenge they face is that even if they lock the versions of dependencies in their project, they cannot publish a lockfile to the package registry, and downstream dependents could not enforce the lockfile for their library. As a result, they choose not to enforce a lockfile in their development environment, as doing so would give them a false sense of security. (“So basically it is better not to commit a lockfile because then you kind of run into these issues faster. And, uh, yeah, you don’t have this false sense of security where you think it’s all locked, but in reality, everybody who uses it doesn’t lock it.”, P4). Following the same reasoning, P8 explain that they prioritize the experience of the users of their library over their own developer experience, and as a result, they do not commit lockfiles.
Similarly, some developers view determinism as a concern of downstream users who depend on their library. (“You could make your end thing deterministic, but as a library author, for me, it’s, that’s not really my problem.” P12).
P9 do not currently commit lockfiles and they mention that they would reconsider their decision based on the updated Cargo guidelines introduced in 2023. They refer to the StackOverflow thread Should Cargo.lock be committed when the crate is both a Rust library and an executable. (“This was the previous recommendation on StackOverflow, but has since changed. I have not revisited this choice, but if I were to start a new project today, I would follow the latest recommendations and defaults”, P9.)
@@ 4.4.2 Slowing down dependency updates
P11 mention that, with Poetry, if a lockfile references a version of a dependency that has been yanked, it results in an error, unlike Cargo. However, if they do not have a lockfile, Poetry will resolve a newer version, without raising an error. They explain the issue with an example. “Now, if you, at 9 a.m. in the morning, let’s say a library pushes a new version. At 9.30, you create a lockfile with that version locked. And at 10 a.m., they delete that version from PyPI. Well, then anybody who syncs using that lockfile hits a big issue. Because that version is no longer available”. At the same time, they acknowledge that this behavior can sometimes be helpful. “it reports it as yanked, and therefore doesn’t install it. So you get an early error, which in some situations is desirable”.
In npm, the current state of project dependencies is referred to as the ideal state. For projects without a lockfile, npm CLI always resolves dependencies to this ideal state at the time of installation. P5 prefer resolving this ideal state rather than relying on a potentially outdated lockfile. “Because you’re always resolving the ideal tree, like at a given moment, that tends to be a better resolution than whatever state your lockfile is accumulated over six months or whatever”. They also believe that continuously resolving the ideal tree helps fixing bugs and vulnerabilities more effectively. “Oftentimes, either you’ll just miss the defect because it will be introduced for a short period of time and then got published with a fix. And so you just won’t even notice it just because you’re always resolving the ideal tree”.
@@ 4.4.3 Human readability challenge
All five npm/pnpm developers mention that they find the lockfiles too long, verbose and hard to read. For example, P1 state “it’s quite wild what you can find inside the package lock. So it’s not really readable”, whereas P2 mention how this verbosity can complicate the code reviews and continuous integration, “Git is kind of overloading and taking up a bunch of memory. And is even having trouble showing them all”. P5 also comment on the human readability of lockfiles, “save us from the insanity that these machine generated lockfiles, which are just, you know, they’re human readable, but they’re not because they’re so big”. Poetry/Pipenv developers raise the same issue where P10 explain, “it’s like a huge JSON file. So it’s kind of annoying to review the diff. And I don’t get a lot of value out of reviewing the diff, necessarily.”. P11 add that they also find it difficult to code review the lockfile, “I will be much more likely to review any changes to a pyproject.toml or, you know, direct dependency file. So I’d say it’s to some degree a weakness of lockfiles that they are so big”.
@@
On the other hand, none of the Cargo or Go developers raise issues related to the readability. Go developer P15 specifically mention they find reviewing go.mod files very easy. “But usually my Go mod files are very like, they do not go over one page of my screen”. P12 who develop in Python and do not use lockfiles also comment that they get a better developer experience with Go, “It’s basically almost the same because go.mod is also kind of very minimalistic, So I guess you get all of the benefits of like, you know, pip lockfile or poetry lockfile or whatever with go without having to think about it.” This confirms the fact that, depending on the lockfile implementation details, such as the content included in lockfiles, developers either effectively use them for manual code reviews or tend to overlook them entirely.
@@
When asked about the information they consider during code reviews of lockfiles, the majority of developers who commit lockfiles mention that they focus on the name, URL, version of the package, and the number of indirect dependencies it includes. The developers do not pay attention to the exact details of indirect dependencies, and trust the package manager to handle them (“I’m peripherally aware that the lockfile has changed. You could say it’s on my mind, but at some point you just have to trust the process, there’s too many dependencies for me to go through everything, for me to do code review on everything.”, P11). A consistent observation we make is that none of the developers pay attention to the checksum of the dependencies in the lockfile. (“I’ve never actually checked checksums”, P7). They do not attempt to read or manually verify the checksums. Instead, they rely on the package manager to automatically verify them. Following the same reasoning, Go developers recommend the Go’s approach of having checksums in a separate file. (“To be honest, I really don’t care where the checksum is stored. It makes it clean to have it separated because I only have to put it into the Go mod file and then the checksum will be generated automatically in the sum file. So that’s like having a look at your Go mod file is much cleaner.”, P15).
@@ 4.4.4 Cache challenge
A common issue faced by all developers is cache invalidation. P8 remark that “Cargo caching mechanisms are buggy”, whereas P1 mention that they have to recreate the lockfile to address cache invalidation problems, “I would say most of the time what people do is they delete the node modules folder, they delete the package lock, and then they retry and install”. Following up on cache invalidation, P2 mention that they occasionally run into operating system specific issues which cannot be solved by simply deleting the cache (“Just working across different operating systems, um, you know, also on macOS. I’ve had issues where deleting, like doing that same process, deleting the node modules folder, created issues with npm and then nothing would install.”). Go developers also face cache related issues with Go proxy when packages get re-released under the same version (“Go proxy already had that basically broken release downloaded and stored in its proxy cache. So people were like opening issues that they had checksum issues”, P15).
@@ 4.4.5 Learning challenge
Another point raised by developers is the learning curve of lockfile related practices across different package managers. P14 compare Gradle and Go noting that they find it easier to learn and understand dependency resolution practices in Go, “it’s very, very simple. It doesn’t have much of a learning curve, such as like Maven or, Gradle”. P12 also comment on this “I think Go is just an easier, it’s just easier. It works better. And, and, you know, it’s a newer language and lessons learned”.
npm developers often mention that the logic behind dependency resolution and lockfile creation processes are unclear (“maybe an understanding of how it constructs the lockfile. Cause it’s still kind of a mystery to me”, P2). Many also comment on the unhelpfulness of error messages when debugging lockfile related issues (“The error message that you get is really cryptic.”, P1). They note that resolving these issues requires an in-depth understanding of the underlying processes, and the debugging itself is extremely time consuming (“Some of these issues I’ve had with it are just so hard to debug until I do some very thorough digging of what’s going on”, P2). P15 compare npm CLI with Go and explain that the minimalism makes the dependency resolution process in Go more transparent and easier to understand. “You just say npm install and like a wall of text goes down and you have a gazillion packages installed and you don’t really know what happens and why those are installed. Compare that to Go, which is, in my opinion, very open and very clear on what dependencies you are using”.
@@ 5.2 Recommendations for package manager developers
6 Threats to Validity
One threat to validity in our study is the selection of keywords used to analyze the implementation of package managers, which may not be exhaustive. We mitigated this threat by combining multiple sources, including both source code and documentation.
The imbalance in the number of projects that use a particular package manager is another potential threat to validity. To minimize bias, we applied the same selection criteria across all package managers and included all projects that satisfied these criteria.
We acknowledge that our dataset and the developers who participated in the qualitative analysis may not represent the entire software engineering practices or the full spectrum of developer perspectives. We also did not interview Gradle developers or Go developers who do not commit the go.sum file, as we considered them outside the norm. This is a threat to validity, as the perspectives of these outliers were not captured. Nevertheless, our results, collected from developers across different ecosystems, remain valid regarding the standard practices within each ecosystem. They provide a broad overview of lockfiles as a general concept across package managers.
@@ 7.1 Dependency resolution and package managers
One of the key responsibilities of a package manager is resolving compatible versions for all the dependencies in a project. Zacchiroli et al.[2] review the usage of dependency resolving algorithms, including specialized constraint solving algorithms as well as ad-hoc dependency tree based algorithms. They conduct a census of dependency resolution mechanisms across 18 package managers, using version locking as one of the criteria for comparison. Pinckney et al.[28] introduce a unified framework built on top of Max-SMT solvers to resolve dependencies more systematically, moving beyond ad-hoc algorithms.
@@
Decan et al.[11] analyze vulnerabilities in the npm dependency network, focusing on how they are discovered and addressed. On the same lines, Liu et al.[20] analyze vulnerability propagation within dependency trees by applying npm-specific dependency resolution rules, rather than relying on simple reachability analysis for indirect dependencies. They recommend managing dependencies with lockfiles. He et al.[14] study the effect of dependency pinning in npm projects. They find that local pinning leads to more security vulnerabilities due to bloated and outdated dependencies. They also suggest that the risk of malicious package updates can be reduced when a small set of core dependencies pin their versions and keep them updated regularly. Rahman et al.[35] confirm the results of He et al.[14], and recommend lockfiles for effective dependency management.
@@ 7.3 Adoption of dependency management tools within development teams
Previous works have also used combined qualitative and quantitative methods to learn about developers’ adoption of dependency management tools. Schorlemmer et al.[37] study the adoption of software signing in four package registries; Maven, PyPI, DockerHub and Huggingface. Kalu et al.[17] conduct a follow-up interview study to understand software signing in practice. They identify that strict signature rules increase the quantity of signatures, and that registry policies have an impact on the developer decision to adopt proper signing. Bifolco et al.[5] assess the accuracy of the GitHub dependency graph in Java and Python projects. They use lockfiles as one source of truth to assess the GitHub dependency graph.
Dependency management bots such as dependabot, renovate and Greenkeeper automate the process of updating dependency specification files and lockfiles when new dependency versions are available. Alfadel et al.[1] and Mohayeji et al.[23] evaluate how developers respond to security updates suggested by Dependabot. Mohayeji et al.[24] also investigate how Dependabot can help mitigate vulnerabilities, noting that it uses lockfiles to create a comprehensive graph of dependencies. Rombaut et al.[31] study whether Greenkeeper reduces developer effort or introduces unnecessary workload. They mention lockfiles as a way to overcome in-range breaking changes, and also briefly mention about the challenge library developers face for sharing lockfiles. Our qualitative analysis confirms that developers do use bots to maintain dependency specification files and lockfiles up to date.